OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP? Use the substitution method to solve the following system of equations.4x - y = 3 7x - 9y = -2 Answers: (1, 1) (6, -3) (6, 1) (1, -3)

OpenStudy (anonymous):

I beg of you to help

OpenStudy (anonymous):

hello.. 4x-y=3 7x - 9y = -2 from the first equation we get : y = 4x -3 substitute it into the second : 7x - 9(4x-3) = -2 7x -36x +27 = -2 -29x = -29 x= 1 substitute x=1 into one of the original equations and get : 4 * 1 -y = 3 y = 1 (1,1)

OpenStudy (anonymous):

oh my god.....you're amazing..................

OpenStudy (anonymous):

how about the 4x4 determinant from that day .. you managed to do it ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh im glad to hear that